OpenStudy (anonymous):

The solution to 2x - 5 = 27 is also a solution of which of the following equations? a.3 + 5x = 58 b.3x - 2 = 31 c.2x + 3 = 35 d.27 - 2x = 5

OpenStudy (anonymous):

@vbroberson @phi

OpenStudy (welshfella):

solving 2x - 5 = 27 add 5 to both sides 2x - 5 + 5 = 27 + 5 2x = 32 divide both sides by 2 x = ?

OpenStudy (welshfella):

can you finish it off/

OpenStudy (anonymous):

idk whats the answer

OpenStudy (welshfella):

2x / 2 = x 32 / 2 = ?

OpenStudy (welshfella):

s = 32/2 = ?

OpenStudy (anonymous):

16 okay but the answer choices are up there

OpenStudy (welshfella):

yea - now what you do is find which choice fits x = 16 for example try the first one:- 3 + 5x = 58 plug in x = 16 3 + 5*16 = 3 + 80 = 83 NOT 58 so it cant be choice 1

OpenStudy (anonymous):

ok......

OpenStudy (anonymous):

now try the others

OpenStudy (welshfella):

for b you want to see if 3*16 - 2 = 31 3*16 - 2 = ?